# Operator: A Protocol for Trustless Delegation Under Uncertainty

### v7

@@ 1 History
These mechanisms relied on the same underlying structure: that if exposing an error was cheaper than committing it, truth became the stable outcome. As Karl Popper observed, knowledge in complex domains depended not on exhaustive specification, but on systematic exposure to falsification [11]. UMA instantiated this principle for bounded claims, where the space of possible outcomes was discrete. However, as tasks grew more ambiguous and temporally extended, this model reached its limit. Operator builds on this work by generalizing adversarial verification to arbitrary tasks. Instead of requiring ex ante clarity or binary outcomes, it enables correctness to be adversarially surfaced, addressing the failure modes of specification and deceptive alignment identified in advanced learning systems [12].
@@ 2 Protocol
Operator enforces correctness by making error economically unprofitable. The protocol is governed by a falsification condition:
@@ 2.1 State Transitions
Figure 1: State transition diagram of the Operator protocol.
@@ 6 Conclusion
We have proposed a system for delegating and verifying arbitrary tasks without relying on specification or trust. By combining collateralized execution, permissionless challenge, and recursive adjudication, Operator ensures that submitting or defending an incorrect result leads to financial loss. The protocol is robust under pseudonymity, resistant to centralized failure, and adaptive to any domain where correctness can be surfaced post hoc. Crucially, as LLM intelligence increases, the spaces where Operator can enforce correctness expand super-linearly: falsification costs $F$ trend toward zero while error detection probability $P_{e}$ approaches one, continuously strengthening the protocol’s effectiveness.

### v9

# A Protocol for Trustless Verification Under Uncertainty

@@ 1 History
These mechanisms relied on the same underlying structure: that if exposing an error was cheaper than committing it, truth became the stable outcome. As Karl Popper observed, knowledge in complex domains depended not on exhaustive specification, but on systematic exposure to falsification [11]. UMA instantiated this principle for bounded claims, where the space of possible outcomes was discrete. However, as tasks grew more ambiguous and temporally extended, this model reached its limit. The protocol builds on this work by generalizing adversarial verification to arbitrary tasks. Instead of requiring ex ante clarity or binary outcomes, it enables correctness to be adversarially surfaced, addressing the failure modes of specification and deceptive alignment identified in advanced learning systems [12].
@@ 2 Protocol
The protocol enforces correctness by making error economically unprofitable. The protocol is governed by a falsification condition:
@@ 2.1 State Transitions
Figure 1: State transition diagram of the protocol.
@@ 6 Conclusion
We have proposed a system for verifying arbitrary tasks without relying on specification or trust. By combining collateralized execution, permissionless challenge, and recursive adjudication, the protocol ensures that submitting or defending an incorrect result leads to financial loss. The protocol is robust under pseudonymity, resistant to centralized failure, and adaptive to any domain where correctness can be surfaced post hoc. Crucially, as LLM intelligence increases, the spaces where the protocol can enforce correctness expand super-linearly: falsification costs $F$ trend toward zero while error detection probability $P_{e}$ approaches one, continuously strengthening the protocol’s effectiveness.